OpenStudy (anonymous):

will medal! simplify. -4.5[2x - (0.5 - 1.4)] a. -9x + 0.9 b. -9x - 0.9 c. -9x - 4.05 d. 9x - 4.05

OpenStudy (anonymous):

@dan815 @heavenfun13 @W.W.G.D @THEGAMER607 @PurplePanda312 @MikeyMaximum @fundays321 @GabbyCakes @mathrulezz @AllTehMaffs @preteenpony @EclipsedStar

OpenStudy (anonymous):

@helpme375

OpenStudy (anonymous):

@Mendicant_Bias

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@SamsungFanBoy

sammixboo (sammixboo):

Are we just simplifying -4.5[2x - (0.5 - 1.4)]

OpenStudy (anonymous):

@sammixboo yeah but i'm stupid and can't figure it out, X~X

sammixboo (sammixboo):

OK so to start simplifying -4.5[2x - (0.5 - 1.4)] we do 0.5 - 1.4 which is?

sammixboo (sammixboo):

And it's OK. You don't understand it, so I am here to help :)

OpenStudy (anonymous):

-0.9?

OpenStudy (anonymous):

brb

sammixboo (sammixboo):

Right! so now we have -4.5(2x - (-9)) Now we must distribute the -4.5, so we multiply -4.5 by EVERYTHING in the parenthesis -4.5 * 2x - (-4.5*-9) So what is -4.5 * -9 and -4.5 * 2x?

sammixboo (sammixboo):

Okie dokie tag me when you are back

OpenStudy (anonymous):

-4.5( 2x - (0.5 - 1.4) ) = -9. x - 4.05

OpenStudy (anonymous):

batman

OpenStudy (anonymous):

back

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

\[-4.5 (2 x-(0.5\, -1.4)) \]\[-4.5 (2 x+0.9)\]\[-9. x-4.05 \]

OpenStudy (anonymous):

@robtobey i don't get it.....X~X

sammixboo (sammixboo):

So continuing where I left at Now we must distribute the -4.5, so we multiply -4.5 by EVERYTHING in the parenthesis -4.5 * 2x - (-4.5*-9) So what is -4.5 * -9 and -4.5 * 2x?

sammixboo (sammixboo):

How -4.5*2x just multiply -4.5 * 2 and then throw on that x

OpenStudy (anonymous):

-4.5 * -9 is 40.5?

sammixboo (sammixboo):

I meant -0.9 sorry

sammixboo (sammixboo):

Aaah, so what is -4.5 * -0.9

OpenStudy (anonymous):

4.05?

sammixboo (sammixboo):

Right, and what is -4.5 * 2x?

sammixboo (sammixboo):

-4.5 * 2x - (-4.5*.-9) -4.5 * 2x - 4.05 To finish simplifying this just multiply -4.5 * 2x

OpenStudy (anonymous):

idk what the "x" in it is so, no clue X~X

sammixboo (sammixboo):

The x is the variable. To multiply two numbers and one has a variable, just multiply the two whole numbers, then throw on the variable. So what is -4.5 * 2

OpenStudy (anonymous):

-9?

sammixboo (sammixboo):

Right, and put the variable x in, so -9x -4.5 * 2x = -9x OK so since that equals that, our answer is -9x - 4.05 !

OpenStudy (anonymous):

4.5 * 0.9 = 45 * 9 * (0.01) = 405 * (0.01) = 4.05

OpenStudy (anonymous):

i don't get it..

sammixboo (sammixboo):

-4.5[2x - (0.5 - 1.4)] -4.5[2x - (-0.9)] -4.5*2x - -4.5(-0.9) -9x - -4.5(-0.9) -9x - 4.05

sammixboo (sammixboo):

Distribution gets a little confusing. I was confused when I first learned it

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

aaanndd i'm doomed...it's 11:55 where i am and i need to get this done before 12:00